Use a suitable identity to get each of the following products.
 (a^2+ b^2 ) (- a ^2+ b^2 )

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a^2+ b^2 ) (- a ^2+ b^2 )
= (b^2+ a^2) (b^2–a^2)
= -a^4+b^4
​Using (a - b)(a + b) = a^2- b^2

Thanks